OpenStudy (anonymous):

integrate sin^5(x) dx. i just need a little point in the right direction!

OpenStudy (anonymous):

\[\sin ^5x~dx=\left( \sin ^2~x \right)^2 \sin xdx= \left( 1-\cos ^2x \right)^2\sin x~dx\] pu t cos x=t

OpenStudy (mathmale):

I assume you mean this: \[\int\limits_{?}^{?}(\sin x)^5 dx\] Is that correct?

OpenStudy (anonymous):

oh yes, that is correct. i am not sure how to type that into the open questions, only when replying.

OpenStudy (anonymous):

@surjithayer that is how i started it actually but i must be doing something wrong after this step

OpenStudy (anonymous):

\[\int\limits \left( 1-\cos ^2x \right)^2\left( \sin x \right)dx\] \[=\int\limits \left( 1-2\cos ^2x+\cos ^4x \right)\sin x~dx=-\cos x+\frac{ 2\cos ^3x }{ 3 }-\frac{ \cos ^5x }{ 5 }+c\]

OpenStudy (anonymous):

or you can solve after substitution cosx=t -sinx dx=dt sin xdx=-dt

OpenStudy (anonymous):

\[\left( 1-t^2 \right)^2=1-2t^2+t4\]

OpenStudy (anonymous):

one brute force method is as follows sin(x)=(ie^(-ix)-ie^(ix))/2 expand (ie^(-ix)-ie^(ix))^5 and integrate

OpenStudy (anonymous):

i ended up learning from you guys and solved it properly! thanks!

OpenStudy (mathmale):

Great! Thank you for acknowledging our help.

OpenStudy (anonymous):

you are welcome.